14) (25 points). Consider the circuit sketched below. If the battery provides an emf of 6.0 V and every resistor has a resistance of 2.0 Ω 2. 7 points. a) What is the equivalent resistance of the resistor network 7 points. b) What are the currents through the resistors Ri and R2. 7 points. c) What are the voltages across the resistors R1 and R2 4 points. d) Which dissipates more power; Ri or R2?
Answer #1

a) Req = R1 + [R2//R3]

= 2 + (2*2)/(2+2)

= 3 ohm

b) I1 = V/Req = 6/3 = 2A

I2 = I1/2 = 1A

c) V1 = I1R1 = 2*2 = 4V

V2 = 1*2 = 2V

d) R1 since it has more current flowing through it.